# OpenWHO: A Document-Level Parallel Corpus for Health Translation in Low-Resource Languages

### v1

@@ 1 Introduction
Our main contributions are:
@@ 2 Related Work
Document-level MT has long been recognised as desirable, as it allows models to leverage broader discourse for improved coherence and accuracy Maruf et al. (2021). Early work with traditional NMT models has demonstrated that document-level context can significantly improve translation quality Miculicich et al. (2018); Post and Junczys-Dowmunt (2024), with Wang et al. (2023) showing that more capable models derive greater benefit from added context. The emergence of LLMs as translation systems has further advanced this paradigm, given their ability to process extended contexts. Koneru et al. (2024) explored contextual translation with Llama2-13B on English-German, finding mixed results, where document-level context sometimes providing no performance gains over sentence-level translation. Karpinska and Iyyer (2023) demonstrated that paragraph-level translation outperforms sentence-level approaches in literary fiction using GPT-3.5, though they have noted that their findings might not generalise to low-resource settings. Yang et al. (2024) fine-tuned LLaMA3-8B for context-aware translation, showing that increasing context window size yields gains, particularly when evaluated with neural metrics. Recent mechanistic analysis by Mohammed and Niculae (2025) revealed that LLMs can be surprisingly “context-insensitive,” with smaller models showing limited ability to effectively utilise available context, a finding that may explain varying performance observed in earlier works.
@@ 4.1 Experimental Setup
We primarily evaluate with ChrF++ Popović (2017), an n-gram based metric which has been shown to correlate better with human judgement than BLEU Papineni et al. (2002) particularly for morphologically rich languages like Kazakh or Georgian. To validate results found with ChrF++, we also evaluate with MetricX-24333google/metricx-24-hybrid-large-v2p6-bfloat16 Juraska et al. (2024) and AutoMQM Fernandes et al. (2023). MetricX is a neural metric which correlates better with human judgement than ChrF++ for high-resource languages. While it has not been evaluated on low-resource languages, it is based on mT5 Xue et al. (2021), which has been pretrained on all languages in our study, aside from Tetun. AutoMQM uses a large language model to characterise translation errors using MQM Lommel et al. (2013). To avoid self-preference bias that may arise from using the same LLM for AutoMQM as that used for translation Wataoka et al. (2025), we run AutoMQM with Kimi K2 Kimi-AI (2025).
@@ 4.2 Results
On the WMT24++ news set, we do not see meaningful accuracy improvements for document-level over sentence-level translation, using either metric (ChrF and MetricX). We alsot see less variation in scores between models on this domain, both when comparing NLLB to LLMs, and when comparing LLMs with each other. Overall, the advantage of document-level translation over sentence-level translation for low-resource MT is not uniform across domains and models.
@@ 6 Conclusion
In this work, we introduced OpenWHO, a high-quality parallel corpus for health MT, with a focus on low-resource languages. Sourced from the World Health Organization’s expert-authored materials, it addresses a gap in evaluation resources and provides a benchmark for future research at the intersection of health and low-resource languages. The dataset strengths include (1) the grounding of its source English text in evidence-based WHO guidance (2) its professional translation into various languages and (3) its availability at both the document and sentence level. However, OpenWHO is language imbalanced (not all courses were translated into all languages), which can limit its comparative value.
@@ 7 Limitations
Our findings rely exclusively on automated metrics (ChrF, MetricX, AutoMQM). While these metrics give a useful signal when they all agree, we have limited ability to resolve differences when they arise. ChrF is a recognised standard for low-resource MT but does not always correlate well with human judgement; MetricX and AutoMQM have not been evaluated on low-resource languages, let alone in the health domain. Overall, more work is needed to determine what is the right metric for low-resource health MT, including a comprehensive human evaluation to validate our findings and gain a more nuanced understanding of translation quality.
@@ 8 Ethics Statement
This work adheres to ethical guidelines for data collection and research in NLP. The OpenWHO corpus was compiled from the World Health Organization’s (WHO) e-learning platform with explicit authorization from the WHO for both data collection and public release. Our work aligns with the WHO’s mission to disseminate health information globally and respects their ownership of the content.
@@ Appendix F Prompts
System: Translate from English to [lang name]. Give only the translation, and no extra commentary, or chattiness. Use the same formatting as the source text to translate, with one sentence per line. Enclose your translation in <result> tags.</result> tags. User: <text to translate>
@@
System: Translate from English to [lang name]. Give only the translation, and no extra commentary, or chattiness. Use the same formatting as the source text to translate, with one sentence per line. Enclose your translation in <result> tags.</result> tags. User: <text to translate>

### v4

@@ 1 Introduction
Our main contributions are  [^1] Code: github.com/raphaelmerx/openwho-code
Dataset: huggingface.co/datasets/raphaelmerx/openwho
@@ 2 Related Work
Document-level MT
Document-level MT has long been recognised as desirable, as it allows models to leverage broader discourse for improved coherence and accuracy Maruf et al. (2021). Early work with traditional NMT models has shown mixed results, with some studies demonstrating that document-level context can significantly improve translation quality Miculicich et al. (2018); Wang et al. (2023); Post and Junczys-Dowmunt (2024), while others questioned whether these improvements stemmed from true contextual understanding, arguing that the context encoder was not modeling discourse but acting as a "noise generator" that improves model robustness, rather than leveraging discourse information Li et al. (2020); Appicharla et al. (2024).
@@
LLMs, given their ability to process extended contexts, are well placed to benefit from document-level context. Koneru et al. (2024) explored contextual translation with Llama2-13B on English-German, finding mixed results, where document-level context sometimes providing no performance gains over sentence-level translation. Karpinska and Iyyer (2023) demonstrated that paragraph-level translation outperforms sentence-level approaches in literary fiction using GPT-3.5, though they have noted that their findings might not generalise to low-resource settings. Yang et al. (2024) fine-tuned LLaMA3-8B for context-aware translation, showing that increasing context window size yields gains, particularly when evaluated with neural metrics. Recent mechanistic analysis by Mohammed and Niculae (2025) revealed that LLMs can be surprisingly “context-insensitive,” with smaller models showing limited ability to effectively utilise available context, a finding that may explain varying performance observed in earlier works.
@@ 4.1 Experimental Setup
We primarily evaluate with ChrF++ Popović (2017), an n-gram based metric which has been shown to correlate better with human judgement than BLEU Papineni et al. (2002) particularly for morphologically rich languages like Kazakh or Georgian. To validate results found with ChrF++, we also evaluate with MetricX-24444google/metricx-24-hybrid-large-v2p6-bfloat16 Juraska et al. (2024) and AutoMQM Fernandes et al. (2023). MetricX is a neural metric which correlates better with human judgement than ChrF++ for high-resource languages. While it has not been evaluated on low-resource languages, it is based on mT5 Xue et al. (2021), which has been pretrained on all languages in our study, aside from Tetun. AutoMQM uses a large language model to characterise translation errors using MQM Lommel et al. (2013). To avoid self-preference bias that may arise from using the same LLM for AutoMQM as that used for translation Wataoka et al. (2025), we run AutoMQM with Kimi K2 Kimi-AI (2025).
@@ 4.2 Results
On the WMT24++ news set, we do not see meaningful accuracy improvements for document-level over sentence-level translation, using either metric (ChrF and MetricX). We also see less variation in scores between models on this domain, both when comparing NLLB to LLMs, and when comparing LLMs with each other. Overall, the advantage of document-level translation over sentence-level translation for low-resource MT is not uniform across domains and models.
@@ 5 Discussion
Our findings, particularly the dependence of context utility on model capability and domain specificity, offer a nuanced picture for where document-level context is most useful, which may explain past work that either did not Li et al. (2020); Appicharla et al. (2024); Koneru et al. (2024) or did Wang et al. (2023); Post and Junczys-Dowmunt (2024); Wu et al. (2024) find added benefits from contextual level translation.
@@ 6 Conclusion
In this work, we introduced OpenWHO, a high-quality parallel corpus for health MT, with a focus on low-resource languages. Sourced from the World Health Organization’s expert-authored materials, it addresses a gap in evaluation resources and provides a benchmark for future research at the intersection of health and low-resource languages. The dataset strengths include (1) the grounding of its source English text in evidence‐based WHO guidance (2) its professional translation into various languages and (3) its availability at both the document and sentence level. However, OpenWHO is language imbalanced (not all courses were translated into all languages), which can limit its comparative value.
@@ 7 Limitations
Our findings rely exclusively on automated metrics (ChrF, MetricX, AutoMQM). While these metrics give a useful signal when they all agree, we have limited ability to resolve differences when they arise. ChrF is a recognised standard for low-resource MT but may not always correlate well with human judgement Wang et al. (2024); MetricX and AutoMQM have not been evaluated on low-resource languages, let alone in the health domain. Overall, more work is needed to determine what is the right metric for low-resource health MT, including a comprehensive human evaluation to validate our findings and gain a more nuanced understanding of translation quality.
@@ 8 Ethics Statement
This work adheres to ethical guidelines for data collection and research in natural language processing. The OpenWHO corpus was compiled from the WHO’s e-learning platform with explicit authorization from the WHO for both data collection and public release. Our work aligns with the WHO’s mission to disseminate health information globally and respects their ownership of the content.
@@ Appendix F Prompts
System: Translate from English to [lang name]. Give only the translation, and no extra commentary, or chattiness. Use the same formatting as the source text to translate, with one sentence per line. Enclose your translation in <result></result> tags. User: <text to translate>
@@
System: Translate from English to [lang name]. Give only the translation, and no extra commentary, or chattiness. Use the same formatting as the source text to translate, with one sentence per line. Enclose your translation in <result></result> tags. User: <text to translate>